# FLUID: Training-Free Face De-identification via Latent Identity Substitution

### v1

@@ Conclusion
In organic chemistry, and especially in combinatorial chemistry, extensive combinations of reagents are used to examine many different types of reaction results. This adds up to our metaphor because such chemical experiments typically do not require time-intensive training-like processes to test the reagent combinations. Similarly, we modify the behavior of the attribute reagent loss by partially fixing components of the source attribute distribution. While our method originally targeted the full distribution of the source image—thereby preserving overall attributes—we now assign varying probability values to specific attributes, ranging from 0.1 to 0.9, to explore targeted manipulations along with de-identification. Results of changing {‘Male’, ‘Narrow Eyes’} and {‘Smile’, ‘Mouth Slightly Open’} are shown in Figure 12. When attributes ‘Smile’ and ‘Mouth Slightly Open’ are assigned low probabilities, the de-identified faces appear neutral. A high probability for ‘Mouth Slightly Open’ and a low one for ‘Smile’ produce an emotionless face with an open mouth. Conversely, assigning a high probability to ‘Smile’ and a low one to ‘Mouth Slightly Open’ generates a closed-mouth smile. High probabilities for both yield a grin. Adjusting the probabilities of ‘Male’ and ‘Narrow Eyes’ also produces interpolated effects. This supports our claim that customizable de-identification can be realized training-free, just like conducting combinatorial chemistry experiments.
Failure Examples and Dataset Bias

### v2

@@ 6.5. Visualizing the Identity-Utility Trade-off
We modify the behavior of the attribute loss by enabling and partially fixing components of the entire 40 attribute distribution. While our method originally targeted only two attributes from the full distribution of the source image—thereby preserving gender and age—we now assign varying probability values to specific attributes, ranging from 0.1 to 0.9, to explore targeted manipulations along with de-identification. Results of changing {‘Male’, ‘Narrow Eyes’} and {‘Smile’, ‘Mouth Slightly Open’} are shown in Figure 10. When attributes ‘Smile’ and ‘Mouth Slightly Open’ are assigned low probabilities, the de-identified faces appear neutral. A high probability for ‘Mouth Slightly Open’ and a low one for ‘Smile’ produce an emotionless face with an open mouth. Conversely, assigning a high probability to ‘Smile’ and a low one to ‘Mouth Slightly Open’ generates a closed-mouth smile. High probabilities for both yield a grin. Adjusting the probabilities of ‘Male’ and ‘Narrow Eyes’ also produces interpolated effects. This supports our claim that customizable de-identification can be realized without training the generative model.
Figure 11. Failure examples and biases from out-of-domain FFHQ edit results.
6.7. Failure Examples and Dataset Bias
@@
7. Limitations and Societal Impact
While our method excels in identity suppression, attribute preservation, and visual quality, it has some limitations. Its iterative latent-space optimization hinders real-time use in resource-limited settings. Performance may drop on highly out-of-distribution data, given its design assumptions and reliance on CelebA-HQ-style priors. Also it may not reliably preserve high-frequency details (e.g., text, glasses, occlusions). Finally, white-box attack against angular distance-based FR models such as MagFace with a more sophisticated design will result in higher identity-utility balances.
8. Conclusion